# Efficient Algorithms for Low Tubal Rank Tensor Approximation with Applications to Image Compression, Super-Resolution and Deep Learning

### v1

@@ Abstract
In this paper we propose efficient randomized fixed-precision techniques for low tubal rank approximation of tensors. The proposed methods are faster and more efficient than the existing fixed-precision algorithms for approximating the truncated tensor SVD (T-SVD). Besides, there are a few works on randomized single-pass algorithms for computing low tubal rank approximation of tensors, none of them experimentally reports the robustness of such algorithms for low-rank approximation of real-world data tensors e.g., images and videos. The current single-pass algorithms for tensors are generalizations of those developed for matrices to tensors. However, the single-pass randomized algorithms for matrices have been recently improved and stabilized. Motivated by this progress, in this paper, we also generalize them to the tensor case based on the tubal product (T-product). We conduct extensive simulations to study the robustness of them compared with the existing single-pass randomized algorithms. In particular, we experimentally found that single-pass algorithms with the sketching parameters of equal sizes usually lead to ill-conditioned tensor least-squares problems and inaccurate results. It is experimentally shown that our proposed single-pass algorithms are robust in this sense. Numerical results demonstrate that under the same conditions (setting the same hyper-parameters), our proposed algorithms provide better performance. Three applications to image compression, super-resolution problem and deep learning are also presented.
@@ 1 Introduction
Tensor SVD (T-SVD) is a widely used tensor decomposition, which has been applied in many applications such as image/video completion [28], face recognition [29], compression [30] and so on. It decomposes a tensor as a T-product (to be defined in Section 2) [31] of three tensors similar to the classical SVD for matrices. To deal with large-scale tensors, several randomized algorithms have been proposed to compute a low tubal rank T-SVD [25, 32, 33]. Recently, we proposed a randomized pass-efficient algorithm in [33] for the T-SVD with an application to image/video completion. However, our algorithm was applicable only for $v\geq 2$ number of passes over the underlying data tensor. It is known that in some applications, we may afford only a one pass and due to this issue, in this paper we would like to propose two efficient single-pass algorithms ($v=1$). To the best of our knowledge, the only paper, which proposes single-pass randomized algorithms for the T-SVD is [32]. The proposed algorithms in [32] are not optimized in several ways. First of all, the authors do not use the fast T-product and t-QR decomposition introduced in [34]. Second, they are generalizations of those proposed in [35] for matrices to tensors, which have been recently modified and stabilised in [36]. It is worth noting that a single-pass algorithm based on sampling lateral and horizontal slices is also developed in [37], see a graphical illustration on this model in Figure 1. In this model, the intersection tensor of the selected lateral and horizontal slices is considered and its Moore-pseudoinverse of it is set as the middle tensor $\underline{\bf U}$ in the CUR approximation. This algorithm does not provide satisfying results when the number of frontal and lateral slices are the same due to the same conditioning properties. We report this in the numerical results and compare it with other single-pass randomized algorithms. To overcome this issue, we present several single-pass algorithms for the computation of the T-SVD and extensively compare their performance with the baselines. In particular, we empirically show that our proposed single-pass algorithms are robust in this sense. Moreover, we apply of proposed single-pass randomized algorithms to the image-super resolution and object detection tasks as important computer vision problems. To best of our knowledge, this paper is the first unitizing single-pass algorithms for such computer vision applications.
@@
We also study the randomized fixed-precision algorithms for computing a low tubal rank approximation of third order tensors. Such algorithms are of interest when a given estimation of the tubal rank is not available and for a given approximation error bound, we have to estimate an appropriate tubal rank and the corresponding low tubal rank approximation. We have recently proposed a fixed-precision algorithm in [38]. Inspired by the interesting results reported in [39], we use these techniques to further improve our randomized fixed-precision algorithm to compute a low tubal rank approximation of third order tensors. The extensive simulations provided in this paper, demonstrate that the new proposed randomized fixed-precision algorithm is faster and more efficient.
@@
    Conducting extensive experimental results on synthetic and real-world data with applications to image/video compression, image super-resolution and deep learning. To the best of our knowledge this is the first paper using single-pass algorithms for tensor completion and image super-resolution.
@@ 2 Background
where $\widehat{{\underline{\bf X}}}(:,:,i)$ is the $i$-th frontal slice of the tensor $\widehat{{\underline{\bf X}}}={\rm fft}({\underline{\bf X}},[],3)$, which calculates the fast Fourier transform of all tubes of the tensor $\underline{\bf X}$, see [34, 25]. Algorithm 2 summarizes the process of T-product of two tensors.
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
The MP pseudoinverse of a tensor can also be computed in the Fourier domain and this is shown in Algorithm 2. The inverse of a tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is denoted by $\underline{\bf X}^{I_{1}\times I_{1}\times I_{3}}$ and is a special case of the MP for which we have ${\underline{\mathbf{X}}^{-1}}*\underline{\mathbf{X}}=\underline{\mathbf{X}}*{\underline{\mathbf{X}}^{-1}}=\underline{\mathbf{I}}$.
@@
(f-diagonal tensor) If all frontal slices of a tensor are diagonal then the tensor is called an f-diagonal tensor.
@@ 1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
3 Generalization of standard matrix decomposition to tensors through the T-product
The T-product may be used to generalize standard matrix decompositions, such as QR, LU, eigenvalue decompositions and SVD, to tenors in a straightforward manner. Given a tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},$ the tensor QR (T-QR) represents a tensor in the form $\underline{\bf X}=\underline{\bf Q}*\underline{\bf R}$ and can be computed through Algorithm 3. We use the notation “orth” for the operator to orthonomalize a given tensor (getting the $\underline{\bf Q}$ part of the T-QR of a given tensor).
It is possible to calculate the tensor LU (T-LU), tensor eigenvalue decomposition (T-EIG) and tensor SVD (T-SVD) by slightly modifying Algorithm 3. More specifically, we substitute the SVD, the LU decomposition and the eigenvalue decomposition of frontal slices $\widehat{\underline{\bf X}}(:,:,i),\,i=1,2,\ldots,I_{3},$ for the QR decomposition in line 4 of Algorithm 3. Please note that Algorithm 3 only needs the thin QR of the first $\lceil\frac{I_{3}+1}{2}\rceil$ slices while the original T-QR algorithm developed in [31, 6] involves the QR of all frontal slices. So, it is recommended to utilize this idea and remove the redundant computations. For example, the T-QR decomposition for a symmetric tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ guarantees the following decomposition
@@ 3 Generalization of standard matrix decomposition to tensors through the T-product
where $\underline{\bf D}$ is f-diagonal. The eigenvalue decomposition (3) can be also represented as follows
@@ 3      $\widehat{\underline{\bf D}}(i,i,j)={\rm sqrt}(\widehat{\underline{\bf D}}(i,i,j))$;
It is worth to also note that the T-SVD is a useful tensor decomposition, which represent a tensor as the T-product of three tensors. The first and last tensors are orthogonal while the middle tensor if f-diagonal. Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be a tensor, then the T-SVD of this tensor is represented as follows
@@
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ and $\underline{\bf V}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$ are orthogonal, while $\underline{\bf S}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is f-diagonal. The number of non-zero diagonal tubes of the the tensor $\underline{\bf S}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is called the tubal rank. The truncated T-SVD is defined by truncating the tensors $\underline{\bf U},\,\underline{\bf S}$, and $\underline{\bf S}$. For example a truncated T-SVD of tubal rank $R$ for the tensor $\underline{\bf X}$ is
@@
where $\underline{\bf U}_{R}=\underline{\bf U}(1:R,:,:)\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf V}_{R}=\underline{\bf U}(:,1:R,:)\in\mathbb{R}^{R\times I_{2}\times I_{3}},$ and $\underline{\bf S}_{R}=\underline{\bf U}(1:R,1:R,:)\in\mathbb{R}^{R\times R\times I_{3}}$. Figure 2 illustrates the structure of this decomposition and its truncated version.
@@ 1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
7       $\widehat{\underline{\mathbf{Q}}}\left({:,:,i}\right)={\rm conj}(\widehat{\underline{\mathbf{C}}}\left({:,:,I_{3}-i+2}\right))$;
@@ 4 Proposed randomized single-pass algorithms
In this paper, we focus on the T-SVD and propose new single-pass algorithms. To he best of our knowledge, the only single-pass algorithms proposed for low tubal rank approximation of tensors were proposed in [37] and [32]. The proposed approach in [37] is based on the cross tensor approximation, that is, for a given tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, a low tubal rank approximation computed by sampling some lateral and horizontal slices of the original data tensor $\underline{\bf X}$, see Figure 1 for an illustration. So, if $\underline{\bf C}\in\mathbb{R}^{I_{1}\times L\times I_{3}}$ and $\underline{\bf R}\in\mathbb{R}^{K\times I_{2}\times I_{3}}$ are some selected lateral and horizontal slices [^2] Please note that it is not necessary the number of lateral and frontal slices be of the same size. of the data tensor $\underline{\bf X}$. Then, a low tubal rank approximation is computed as follows
@@
So, this single-pass or 1-view method combines range and co-range sketches in a single pass over the data tensor $\underline{\bf X}$, and then generate a low tubal rank estimate based on information included in the sketches. These algorithms is summarized in Algorithms 5. Consider a streaming setting where the data tensor $\underline{\bf X}$ is never entirely kept in random-access memory but is instead given as a finite stream of linear updates.
@@
As discussed in [36], the single-pass algorithms for matrices usually give unreliable approximations when $L=K$ due to some instability issues in solving badly conditioned least-squares problems. Actually, we experienced the same for tensors and will report them in our simulations. According to (10), we need to solve the linear tensor equation $(\underline{\bf\Omega}_{2}*\underline{\bf Q})*\underline{\bf Y}=\underline{\bf W}.$ The coefficient tensor $\underline{\bf\Omega}_{2}*\underline{\bf Q}\in\mathbb{R}^{L\times K\times I_{3}}$ has $L$ and $K$ number of horizontal and lateral slices, respectively. One way to avoid the badly conditioned coefficient tensor is considering $L>K$ and solving an overdetermined and better-conditioned problem. However, this requires fine-tuning the parameters $L$ and $K$ to get an acceptable approximation, imposing extra costs of carefully identifying the best of such values, otherwise there is a the risk of losing accuracy. On the other hand, as highlighted in [36], one other reason to address the disadvantages of selecting $L=K$ is that $K$ may represent a processing bottleneck in some applications. Given the computing limitations, it would thus be tempting to set $L=K$ in an attempt to obtain as much information as possible. Nevertheless, the requirement to retain both the sketching and random sample matrices places a memory restriction on the 1-view approach. The last comment that needs to be mentioned is that using prior information from the singular values of the matrix at hand in [35], several ways are suggested to select $L$ and $K$, which are of less practical interest when we have no access to such information. It is worth mentioning that in all our proposed single-pass algorithms, the sketches $\underline{\bf Y}_{c}$ or $\underline{\bf Y}_{r}$ are much smaller than the original data tensor, and computing their left singular vectors will not be very expensive compared to only an orthonormal basis for the range sketch. Particularly since, in single-pass techniques, multiplying $\underline{\bf X}$ to random tensors (to discover range and co-range sketches) usually imposes the most computational cost and time.
@@
To overcome this drawback, more stabilized and improved algorithms (see Algorithms 7 and 8 in [36]) were developed in [36] for matrices. Motivated by their efficiency, we generalized them to the tensor case and are summarized in Algorithms 6 and 7. The principal idea is to use the truncated T-SVD for getting an orthonormal basis for the full range of the sketches $\underline{\bf Y}_{r}$ or $\underline{\bf Y}_{c}$ instead of applying the T-QR decomposition. Let us explain this in detail. Instead of considering $\underline{\bf Q}$ as an orthonomal basis for $\underline{\bf Y}$, we compute $[\underline{\bf Q},\sim,\sim]={\rm Truncated\,T\operatorname{-}SVD}(\underline{\bf Y},H)$, where $0\leq H\leq K$. This means that $\underline{\bf Q}\in\mathbb{R}^{I_{1}\times H\times I_{3}}$ contains the leading left singular tensors of $\underline{\bf Y}$, and if $H\leq K$, then coefficient tensor $\underline{\bf\Omega}_{2}*\underline{\bf Q}\in\mathbb{R}^{L\times H\times I_{3}}$ should provide a more conditioned problem, even for $L=K.$
@@
Note that other single-pass methods, such as two-sided randomized SVD (TSR-SVD) [41], can also be extended to the tensor case. The TSR-SVD uses the compressed tensor $\underline{\bf W}=\underline{\bf\Omega}_{2}*\underline{\bf X}*\underline{\bf\Omega}_{1},$ where $\underline{\bf\Omega}_{1}$ and $\underline{\bf\Omega}_{2}$ are standard Gaussian tensors of conforming sizes, to compute a low tubal rank approximation of a tensor. To be more precise, for a data tensor $\underline{\bf X}$, and orthogonal bases $\underline{\bf Q}_{1}$ and $\underline{\bf Q}_{2}$, we have
@@
It is not difficult to see that the middle tensor in (11), can be approximated as follows
@@
Both of these require one to pass over the underlying data matrix.

### v2

@@ Abstract
In this paper we propose efficient randomized fixed-precision techniques for low tubal rank approximation of tensors. The proposed methods are faster and more efficient than the existing fixed-precision algorithms for approximating the truncated tensor SVD (T-SVD). Besides, there are a few works on randomized single-pass algorithms for computing low tubal rank approximation of tensors, none of them experimentally reports the robustness of such algorithms for low-rank approximation of real-world data tensors, e.g., images and videos. The current single-pass algorithms for tensors are generalizations of those developed for matrices to tensors. However, the single-pass randomized algorithms for matrices have been recently improved and stabilized. Motivated by this progress, in this paper, we also generalize them to the tensor case based on the tubal product (T-product). We conduct extensive simulations to study the robustness of them compared with the existing single-pass randomized algorithms. In particular, we experimentally found that single-pass algorithms with sketching parameters of equal sizes usually lead to ill-conditioned tensor least-squares problems and inaccurate results. It is experimentally shown that our proposed single-pass algorithms are robust in this sense. Numerical results demonstrate that under the same conditions (setting the same hyperparameters), our proposed algorithms provide better performance. Three applications to image compression, super-resolution problems, and deep learning are also presented.
@@ 1 Introduction
Tensor SVD (T-SVD) is a widely used tensor decomposition, which has been applied in many applications such as image/video completion [28], face recognition [29], compression [30] and so on. It decomposes a tensor as a T-product (to be defined in Section 2) [31] of three tensors similar to the classical SVD for matrices. To deal with large-scale tensors, several randomized algorithms have been proposed to compute a low tubal rank T-SVD [25, 32, 33]. Recently, we proposed a randomized pass-efficient algorithm in [33] for the T-SVD with an application to image/video completion. However, our algorithm was applicable only for $v\geq 2$ number of passes over the underlying data tensor. It is known that in some applications, we may afford only one pass and due to this issue, in this paper, we would like to propose two efficient single-pass algorithms ($v=1$). To the best of our knowledge, the only paper, which proposes single-pass randomized algorithms for the T-SVD is [32]. The proposed algorithms in [32] are not optimized in several ways. First of all, the authors do not use the fast T-product and t-QR decomposition introduced in [34]. Second, they are generalizations of those proposed in [35] for matrices to tensors, which have been recently modified and stabilized in [36]. It is worth noting that a single-pass algorithm based on sampling lateral and horizontal slices is also developed in [37], see a graphical illustration of this model in Figure 1. In this model, the intersection tensor of the selected lateral and horizontal slices is considered, and its Moore-pseudoinverse is set as the middle tensor $\underline{\bf U}$ in the CUR approximation. This algorithm does not provide satisfying results when the number of frontal and lateral slices are the same due to the conditioning issue. We report this in the numerical results and compare it with other single-pass randomized algorithms. To overcome these issues with the current single-pass algorithms, we present several single-pass algorithms for the computation of the T-SVD and extensively compare their performance with the baselines. In particular, we empirically show that our proposed single-pass algorithms are robust in this sense. Moreover, we apply the proposed single-pass randomized algorithms to the image super-resolution and object detection tasks as important computer vision problems. To the best of our knowledge, this paper is the first to unitize single-pass algorithms for such computer vision applications.
@@
We also study the randomized fixed-precision algorithms for computing a low tubal rank approximation of third order tensors. Such algorithms are of interest when a given estimation of the tubal rank is not available, and for a given approximation error bound, we have to estimate an appropriate tubal rank and the corresponding low tubal rank approximation. We have recently proposed a fixed-precision algorithm in [38]. Inspired by the interesting results reported in [39], we use these techniques to further improve our randomized fixed-precision algorithm to compute a low tubal rank approximation of third order tensors. The extensive simulations provided in this paper, demonstrate that the new proposed randomized fixed-precision algorithm is faster and more efficient.
@@
    Conducting extensive experimental results on synthetic and real-world data with applications to image/video compression, image super-resolution, and deep learning. To the best of our knowledge, this is the first paper using single-pass algorithms for tensor completion and image super-resolution.
@@ 2 Background
where $\widehat{{\underline{\bf X}}}(:,:,i)$ is the $i$-th frontal slice of the tensor $\widehat{{\underline{\bf X}}}={\rm fft}({\underline{\bf X}},[],3)$, which calculates the fast Fourier transform of all tubes of the tensor $\underline{\bf X}$, see [34, 25]. Algorithm 1 summarizes the process of the T-product of two tensors.
@@ 2 $\widehat{\underline{\mathbf{Y}}}={\rm fft}\left({\underline{\mathbf{Y}},[],3}\right)$;
The MP pseudoinverse of a tensor can also be computed in the Fourier domain, and this is shown in Algorithm 2. The inverse of a tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ is denoted by $\underline{\bf X}^{I_{1}\times I_{1}\times I_{3}}$ and is a special case of the MP for which we have ${\underline{\mathbf{X}}^{-1}}*\underline{\mathbf{X}}=\underline{\mathbf{X}}*{\underline{\mathbf{X}}^{-1}}=\underline{\mathbf{I}}$.
@@
(f-diagonal tensor) If all frontal slices of a tensor are diagonal, then the tensor is called an f-diagonal tensor.
@@ 1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
3 Generalization of standard matrix decompositions to tensors through the T-product
The T product may be used to generalize standard matrix decompositions, such as QR, LU, eigenvalue decompositions, and SVD, to tensors in a straightforward manner. Given a tensor, $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}},$ the tensor QR (T-QR) represents a tensor in the form $\underline{\bf X}=\underline{\bf Q}*\underline{\bf R}$ and can be computed through Algorithm 3. The MATLAB command “qr(X,0)”, means the economic QR decomposition of the matrix ${\bf X}$ in Line 3 of the algorithm.
It is possible to calculate the tensor LU (T-LU), tensor eigenvalue decomposition (T-EIG) and tensor SVD (T-SVD) by slightly modifying Algorithm 3. More specifically, we substitute the SVD, the LU decomposition, and the eigenvalue decomposition of frontal slices $\widehat{\underline{\bf X}}(:,:,i),\,i=1,2,\ldots,I_{3},$ with the QR in line 4 of Algorithm 3. Please note that Algorithm 3 only needs the thin QR of the first $\lceil\frac{I_{3}+1}{2}\rceil$ slices, while the original T-QR algorithm developed in [31, 6] involves the QR of all frontal slices. So, it is recommended to utilize this idea and remove the redundant computations. For example, the T-EIG decomposition for a symmetric tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ guarantees the following decomposition
@@ 3 Generalization of standard matrix decompositions to tensors through the T-product
where $\underline{\bf D}$ is f-diagonal. The eigenvalue decomposition (3) can also be represented as follows,
@@ 3      $\widehat{\underline{\bf D}}(i,i,j)={\rm sqrt}(\widehat{\underline{\bf D}}(i,i,j))$;
It is worth also noting that the T-SVD is a useful tensor decomposition, representing a tensor as the T-product of three tensors. The first and last tensors are orthogonal, while the middle tensor if f-diagonal. Let $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ be a tensor; then the T-SVD of this tensor is represented as follows,
@@
where $\underline{\bf U}\in\mathbb{R}^{I_{1}\times I_{1}\times I_{3}}$ and $\underline{\bf V}\in\mathbb{R}^{I_{2}\times I_{2}\times I_{3}}$ are orthogonal, while $\underline{\bf S}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is f-diagonal. The number of non-zero diagonal tubes of the tensor $\underline{\bf S}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ is called the tubal rank. The truncated T-SVD is defined by truncating the tensors $\underline{\bf U},\,\underline{\bf S}$, and $\underline{\bf S}$. For example a truncated T-SVD of tubal rank $R$ for the tensor $\underline{\bf X}$ is
@@
where $\underline{\bf U}_{R}=\underline{\bf U}(:,1:R,:)\in\mathbb{R}^{I_{1}\times R\times I_{3}},\,\underline{\bf V}_{R}=\underline{\bf V}(:,1:R,:)\in\mathbb{R}^{R\times I_{2}\times I_{3}},$ and $\underline{\bf S}_{R}=\underline{\bf S}(1:R,1:R,:)\in\mathbb{R}^{R\times R\times I_{3}}$. Figure 2 illustrates the structure of this decomposition and its truncated version.
@@ 1 $\widehat{\underline{\mathbf{X}}}={\rm fft}\left({\underline{\mathbf{X}},[],3}\right)$;
7       $\widehat{\underline{\mathbf{Q}}}\left({:,:,i}\right)={\rm conj}(\widehat{\underline{\mathbf{Q}}}\left({:,:,I_{3}-i+2}\right))$;
@@ 4 Proposed randomized single-pass algorithms
In this paper, we focus on the T-SVD and propose new single-pass algorithms. To the best of our knowledge, the only single-pass algorithms proposed for low tubal rank approximation of tensors were proposed in [37] and [32]. The proposed approach in [37] is based on the cross tensor approximation; that is, for a given tensor $\underline{\bf X}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, a low tubal rank approximation is computed by sampling some lateral and horizontal slices of the original data tensor $\underline{\bf X}$, see Figure 1 for an illustration. So, if $\underline{\bf C}\in\mathbb{R}^{I_{1}\times L\times I_{3}}$ and $\underline{\bf R}\in\mathbb{R}^{K\times I_{2}\times I_{3}}$ are some selected lateral and horizontal slices [^2] Please note that it is not necessary the number of lateral and frontal slices be of the same size. of the data tensor $\underline{\bf X}$. Then, a low tubal rank approximation is computed as follows,
@@
So, this single-pass or 1-view method combines range and co-range sketches in a single pass over the data tensor $\underline{\bf X}$, and then generates a low tubal rank estimate based on information included in the sketches. These algorithms is summarized in Algorithm 5. Consider a streaming setting where the data tensor $\underline{\bf X}$ is never entirely kept in random-access memory but is instead given as a finite stream of linear updates.
@@
As discussed in [36], the single-pass algorithms for matrices usually give unreliable approximations when $L=K$ due to some instability issues in solving badly conditioned least-squares problems. Actually, we experienced the same for tensors and will report them in our simulations. According to (10), we need to solve the linear tensor equation $(\underline{\bf\Omega}_{2}*\underline{\bf Q})*\underline{\bf Y}=\underline{\bf W}.$ The coefficient tensor $\underline{\bf\Omega}_{2}*\underline{\bf Q}\in\mathbb{R}^{L\times K\times I_{3}}$ has $L$ and $K$ number of horizontal and lateral slices, respectively. One way to avoid the badly conditioned coefficient tensor is to consider $L>K$ and solve an overdetermined and better-conditioned problem. However, this requires fine-tuning the parameters, $L$ and $K$ to get an acceptable approximation, imposing extra costs of carefully identifying the best of such values; otherwise, there is the risk of losing accuracy. On the other hand, as highlighted in [36], one other reason to address the disadvantages of selecting $L=K$ is that $K$ may represent a processing bottleneck in some applications. Given the computing limitations, it would thus be tempting to set $L=K$ in an attempt to obtain as much information as possible. Nevertheless, the requirement to retain both the sketching and random sample matrices places a memory restriction on the 1-view approach. The last comment that needs to be mentioned is that using prior information from the singular values of the matrix at hand in [35], several ways are suggested to select $L$ and $K$, which are of less practical interest when we have no access to such information. It is worth mentioning that in all our proposed single-pass algorithms, the sketches $\underline{\bf Y}_{c}$ or $\underline{\bf Y}_{r}$ are much smaller than the original data tensor, and computing their left singular vectors will not be very expensive compared to only an orthonormal basis for the range sketch. Particularly since, in single-pass techniques, multiplying $\underline{\bf X}$ with random tensors (to discover range and co-range sketches) usually imposes the most computational cost and time.
@@
To overcome this drawback, more stabilized and improved algorithms (see Algorithms 7 and 8 in [36]) were developed in [36] for matrices. Motivated by their efficiency, we generalized them to the tensor case and are summarized in Algorithms 6 and 7. The principal idea is to use the truncated T-SVD for getting an orthonormal basis for the full range of the sketches $\underline{\bf Y}_{r}$ or $\underline{\bf Y}_{c}$ instead of applying the T-QR decomposition. Let us explain this in detail. Instead of considering $\underline{\bf Q}$ as an orthonormal basis for $\underline{\bf Y}$, we compute $[\underline{\bf Q},\sim,\sim]={\rm Truncated\,T\operatorname{-}SVD}(\underline{\bf Y},H)$, where $0\leq H\leq K$. This means that $\underline{\bf Q}\in\mathbb{R}^{I_{1}\times H\times I_{3}}$ contains the leading left singular tensors of $\underline{\bf Y}$, and if $H\leq K$, then the coefficient tensor $\underline{\bf\Omega}_{2}*\underline{\bf Q}\in\mathbb{R}^{L\times H\times I_{3}}$ should provide a more conditioned problem, even for $L=K.$
@@
Note that other single-pass methods, such as two-sided randomized SVD (TSR-SVD) [41], can also be extended to the tensor case. The TSR-SVD uses the compressed tensor, $\underline{\bf W}=\underline{\bf\Omega}_{2}*\underline{\bf X}*\underline{\bf\Omega}_{1},$ where $\underline{\bf\Omega}_{1}$ and $\underline{\bf\Omega}_{2}$ are standard Gaussian tensors of conforming sizes, to compute a low tubal rank approximation of a tensor. To be more precise, for a data tensor $\underline{\bf X}$, and orthogonal bases $\underline{\bf Q}_{1}$ and $\underline{\bf Q}_{2}$, we have
@@
It is not difficult to see that the middle tensor in (11), can be approximated as follows,
@@
Both of these require one to pass over the underlying data tensor.